# iCost: A Novel Instance Complexity Based Cost-Sensitive Learning Framework for Imbalanced Classification

### v1

@@ IV-D Performance Comparison
We compared the performance of the proposed approach with the standard CS learning technique to evaluate the differences. The results have also been compared with popular sampling techniques used in imbalanced learning. These include oversampling approaches such as SMOTE, ADASYN, Borderline-SMOTE (BL-SMOTE), and Random oversampling (ROS); undersampling techniques such as Random undersampling (RUS), Edited nearest neighbor (ENN), and Neighborhood cleaning (NC); hybrid sampling technique SMOTE-Tomek. Details of the performance measures from these approaches have been reported in the following section. The imblearn library was utilized to implement these sampling techniques using default parameter settings.
V Results and Discussion
@@
In this section, we present and discuss the results we have obtained during the experiment. We have measured the performance of 3 different classifiers on 66 imbalanced datasets using 8 different metrics. All these measures cannot be included here due to space constraints. They are provided in separate supplementary files. We mainly considered the MCC, ROC-AUC, G-mean, and F1-score for comparison.
V-A Performance from traditional approaches
@@
Standard classification algorithms do not perform well on imbalanced datasets. The performance gets even worse with higher imbalances and class overlapping [34]. In many cases, the g-mean score was found to be 0 (sensitivity=0, specificity=100), indicating a clear bias towards the majority class. Cost-sensitive learning (CSL) provides a significant improvement in performance as can be observed from Fig 3. Fig. 3 shows the average G-mean scores obtained on 66 datasets. Among the classifiers, CS-SVM provided the highest G-mean score. DT is found to be less sensitive to CS approaches. Improvement in performance was the maximum in the LR classifier. The average results from other metrics are provided in Table IV, Table V, and Table VI.
@@
V-B Performance comparison of the proposed iCost algorithm with the standard CS approach
@@
Figure 4: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the LR classifier on 66 datasets.
@@
Figure 5: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the SVM classifier on 66 datasets.
@@
Figure 6: Change in MCC score from the iCost algorithm as compared to traditional CS approach for the DT classifier on 66 datasets.
@@
The changes in the MCC score obtained from the proposed algorithm as compared to the traditional CS approach are presented in Fig. 4, Fig. 5, and Fig. 6 for the LR, SVM, and DT classifiers, respectively. As can be observed from Fig. 4, a significant improvement is noticeable in most datasets for the LR classifier. For a few datasets, there was no change in performance and for only one dataset, the performance dropped slightly. As for the SVM classifier, performance also improved in most datasets. However, the increment was lower than the LR classifier. In several of the datasets, the performance remained unchanged. This happened mostly in highly imbalanced cases (the datasets are sorted based on IR in ascending order). In such cases, the number of minority class samples is usually limited and they are surrounded by instances of the opposite class, leaving very few pure and safe samples. As a result, almost all the samples are border samples which are weighted equally and the algorithm works the same as the standard CS approach. In the case of the DT classifier, there was a deterioration in performance in some of the datasets. However, in most cases, it was almost negligible. As for the other datasets, there was a decent improvement in performance.
@@
Overall, the most improvement was noticeable in the LR classifier (3.3% on average per dataset). For the other two classifiers, the improvement was around 2% on average. MCC is a highly reliable performance metric, and an improvement in MCC indicates that our proposed method effectively reduces the number of misclassifications. Fig. 7 demonstrates the other performance measures (average) obtained on 66 imbalanced datasets for the LR classifier. As can be observed from the figure, there is an increase in performance from the proposed algorithm over the traditional CS approach for almost all the measures. Only the sensitivity score dropped marginally. The number of minority class samples available in the datasets is usually very small. As a result, only a few misclassifications lead to a large change in the sensitivity score. However, in terms of all four composite metrics, the proposed algorithm achieved better results compared to the traditional CS approach. Similar improvements are noticeable for the SVM classifier as well (Table V.
Figure 7: Average performance measures on 66 datasets for the LR classifier.
@@
V-C Performance comparison of the proposed algorithm with other sampling techniques
@@
We also compared the performance with some of the popular sampling approaches. Data resampling techniques undertake a completely different path in addressing the class imbalance and they have been widely applied in the imbalanced domain. The average of performance measures obtained from these approaches on the 66 imbalanced datasets are provided in Table IV, Table V, and Table VI for the LR, SVM, and DT classifiers, respectively.
@@
As can be observed from the tables, the proposed framework produced better prediction performance compared to all the other approaches in terms of precision, ROC-AUC, G-mean, MCC, and F1-score. When the data was resampled using SMOTE, the most popular sampling technique, G-mean scores of 77.62% and 78.45% were obtained for the LR and SVM classifiers, respectively. Our proposed approach outperformed this well-established method by a large margin producing G-mean scores of 81.03% and 82.8%, respectively. A significant improvement in performance was noticeable compared to other approaches as well.
@@
For both cases, the standard classifiers provided the highest specificity score due to the bias towards the majority class. For the SVM classifier, RUS provided the highest sensitivity score. However, RUS also attained the lowest specificity score. In RUS, a large number of majority-class samples are removed from the data to attain balance. Consequently, this approach provided the best performance for the minority class at the cost of the lowest performance for the majority class. This makes the prediction framework quite unreliable. A higher sensitivity score does not really indicate that the approach is actually performing well. MCC score manifests a more clear picture and the RUS approach produced a MCC score of 48.15%. Compared to that, the iCost algorithm achieved a much higher MCC score of 60.02%.
@@
TABLE IV: Performance measures obtained from different approaches for the LR classifier
Metrics LR SMOTE ADASYN BL-SMOTE ROS RUS ENN NC SMOTE_Tomek CS-LR iCost (Proposed)
@@
TABLE V: Performance measures obtained from different approaches for the SVM classifier
Metrics SVM SMOTE ADASYN BL_SMOTE ROS RUS ENN NC SMOTE_Tomek CS-SVM iCost(Proposed)
@@
TABLE VI: Performance measures obtained from different approaches for the DT classifier
Metrics DT SMOTE ADASYN BL_SMOTE ROS RUS ENN NC SMOTE_Tomek CS-DT iCost(Proposed)
@@
VI Conclusion
@@
In this study, a modified Cost-sensitive learning framework is proposed where data difficulty factors are taken into consideration while penalizing the instances. In the traditional framework, all the instances are weighted equally. This is impractical and can bias the prediction toward the minority-class, leading to an increased amount of false positives. Uniform instance weighting overfits the data by unusually deforming the decision boundary. This does not fare well during testing, resulting in a higher number of misclassifications. Class overlapping is identified as the most crucial difficulty factor that impairs prediction performance. To alleviate the issue, we weigh the minority-class samples in the overlapping region more strongly compared to the ones in the non-overlapping region. Carefully assigning higher weights to the hard-to-learn examples while reducing the weights of the other provides a more plausible weighting mechanism, resulting in fewer misclassifications. We have tested our algorithm on 66 imbalanced datasets using 3 different classifiers and observed an increment in performance in most cases. The proposed algorithm is computationally light similar to traditional CS approaches. The modification improves the CS framework by introducing some reasonable adjustments to it.
@@
Only binary imbalanced classification scenarios have been considered in this study. The concept can be extended for multiclass scenarios as well. We plan to explore that in future works. While we worked with three different classifiers in this study, the proposed algorithm can be used with other classification algorithms such as Random Forest or XGBoost. The default values for the proposed approach are empirically set. Further research is required in this regard to understand how different data difficulty factors are related to the cost-factor values. We measured the instance complexity based on nearest neighbors. However, there are other data complexity measures such as local sets [35] which can also be considered. We plan to incorporate that into our framework in the future.
Supplementary files and codes are available in this repository: (could not be released during submission due to IEEE’s blind review policy).

### v2

@@ 4.4 Performance Comparison
There are many other techniques available that are used in imbalanced learning. Data resampling is a completely different paradigm in imbalanced data handling. As such, comparing performance with state-of-the-art sampling techniques is irrelevant since that is not the primary objective of the proposed algorithm. Nevertheless, we compare our results with some of the popular sampling techniques. These include oversampling approaches such as SMOTE [31], ADASYN [54], Borderline-SMOTE (BLS) [55], and Random oversampling (ROS); undersampling techniques such as Random undersampling (RUS), Edited nearest neighbor (ENN), and Neighborhood cleaning (NC); hybrid sampling technique SMOTE-Tomek. Details of the performance measures from these approaches have been reported in the following section. The imblearn library was utilized to implement these sampling techniques using the default parameter settings [56].
5 Results and Discussion
@@
In this section, we present and discuss the results we have obtained during the experiment. We have measured the performance of 4 different classifiers on 65 binary imbalanced datasets and 10 multiclass imbalanced datasets using 8 different metrics. All these measures cannot be included here due to space constraints. They are provided in separate supplementary files. We mainly considered the MCC, ROC-AUC, G-mean, and F1-score for comparison. The results from the ’neighborhood’ categorization and type = ’ins’ are elaborately presented here and have also been applied to the multiclass scenario.
5.1 Performance from traditional approaches
@@
Standard classification algorithms do not perform well on imbalanced datasets. The performance gets even worse with higher imbalances and class overlapping [30]. In many cases, the g-mean score was found to be 0 (sensitivity=0, specificity=100), indicating a clear bias towards the majority class. CSL provides a good improvement in performance as can be observed from Fig 6. It shows the average G-mean scores obtained on 65 datasets. Among the classifiers, CS-SVM provided the highest G-mean score. Improvement in performance was the maximum in the LR classifier. DT is found to be less sensitive to CS approaches. For the RF classifier, the average performance actually dropped a little when CSL was applied.
One thing to note here is that this is the average g-mean score obtained on the 65 datasets we tested. It does not portray the individual cases. In many of them, the performance improved from the CS-RF or CS-DT approaches. While in many others the performance dropped and on average, there was a decline in the overall g-mean score. The average results from other metrics are provided in Table 5, Table 6, Table 7, and Table 8.
@@
5.2 Performance comparison of the proposed iCost algorithm with the standard CS approach
@@
Figure 7: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the LR classifier on 65 datasets.
@@
Figure 8: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the SVM classifier on 65 datasets.
@@
Figure 9: Change in MCC score from the iCost algorithm as compared to traditional CS approach for the DT classifier on 65 datasets.
Figure 10: Change in MCC score from the iCost algorithm as compared to traditional CS approach for the RF classifier on 65 datasets.
The changes in the MCC score obtained from the proposed algorithm as compared to the traditional CS approach are presented in Fig. 7, Fig. 8, Fig. 9, and Fig. 10 for the LR, SVM, DT, and RF classifiers, respectively. The datasets are sorted based on IR in ascending order on the x-axis. Other measures of performance on individual datasets are provided as supplementary files in the GitHub repository.
@@
As can be observed from Fig. 7, a significant improvement is noticeable in most datasets for the LR classifier. For a few datasets, there was no change in performance and for only one dataset, the performance dropped slightly. As for the SVM classifier, performance also improved in most datasets. However, the increment was lower than the LR classifier. In several of the datasets, the performance remained unchanged. This happened mostly in highly imbalanced cases. In such cases, the number of minority class samples is usually limited and they are surrounded by instances of the opposite class, leaving very few pure and safe samples. As a result, almost all the samples are border samples which are weighted equally and the algorithm works the same as the standard CS approach. In the case of the DT classifier, there was a deterioration in performance in some of the datasets. However, in most cases, it was almost negligible. As for the other datasets, there was a decent improvement in performance. For the RF classifier, there was an overall improvement in most cases. Performance dropped marginally for a couple of datasets.
@@
Overall, the most improvement was noticeable in the LR classifier (3.3% on average per dataset). For the other three classifiers, the improvement was around 2% on average. MCC is a highly reliable performance metric, and an improvement in MCC indicates that our proposed method effectively reduces the number of misclassifications. Fig. 11 demonstrates the other performance measures (average) obtained on 65 imbalanced datasets for the LR classifier. As can be observed from the figure, there is an increase in performance from the proposed algorithm over the traditional CS approach for almost all the measures. Only the sensitivity score dropped marginally. The number of minority class samples available in the datasets is usually very small. As a result, only a few misclassifications lead to a large change in the sensitivity score. However, in terms of all four composite metrics, the proposed algorithm achieved better results compared to the traditional CS approach. For the RF classifier, both sensitivity and specificity scores improved by employing the proposed framework. Similar improvements are noticeable for other classifiers.
Figure 11: Average performance measures on 65 datasets for the LR classifier.
@@
5.3 Performance comparison of the proposed algorithm with other sampling techniques
@@
The performance was also compared with some of the popular sampling approaches. Data resampling techniques undertake a completely different path in addressing the class imbalance and they have been widely applied in imbalanced learning tasks. The average of performance measures obtained from these approaches on the 65 imbalanced datasets are provided in Table 5, Table 6, Table 7, and Table 8 for the LR, SVM, DT, and RF classifiers, respectively.
@@
As can be observed from the tables, the proposed framework usually produced better prediction performance compared to the other approaches in terms of precision, ROC-AUC, G-mean, MCC, and F1-score. When the data was resampled using SMOTE, the most popular sampling technique, G-mean scores of 77.62% and 78.45% were obtained for the LR and SVM classifiers, respectively. Our proposed approach outperformed this well-established method by producing G-mean scores of 81.03% and 82.8%, respectively. A good improvement in performance was noticeable compared to other approaches as well.
For the RF classifier, however, the sampling techniques performed quite well compared to the CSL approach. While the iCost algorithm outperformed the traditional CS-RF approach, it could not outdo other sampling approaches. RF is an ensemble method that works by creating bootstrap subsets. In each bootstrap subset, the IR is different than the original dataset. The CS algorithms do not capture that as it uses the overall IR. That is why the performance gets constantly lower compared to the sampling techniques. We plan to address this issue in the future.
@@
For all cases, the standard classifiers provided the highest specificity score due to the bias towards the majority class. For the SVM classifier, RUS provided the highest sensitivity score. However, RUS also attained the lowest specificity score. In RUS, a large number of majority-class samples are removed from the data to attain balance. Consequently, this approach provided the best performance for the minority class at the cost of the lowest performance for the majority class. This makes the prediction framework quite unreliable. A higher sensitivity score does not really indicate that the approach is actually performing well. Metrics such as ROC-AUC or MCC score manifests a more clear picture and the RUS approach produced a MCC score of 48.15%. Compared to that, the iCost algorithm achieved a much higher MCC score of 60.02%.
@@
Table 5: Performance measures obtained from different approaches for the LR classifier
Metrics LR SMOTE ADASYN BLS ROS RUS ENN NC S_Tomek CS-LR iCost (Proposed)
@@
Table 6: Performance measures obtained from different approaches for the SVM classifier
Metrics SVM SMOTE ADASYN BLS ROS RUS ENN NC S_Tomek CS-SVM iCost(Proposed)
@@
Table 7: Performance measures obtained from different approaches for the DT classifier
Metrics DT SMOTE ADASYN BLS ROS RUS ENN NC S_Tomek CS-DT iCost(Proposed)
@@
Table 8: Performance measures obtained from different approaches for the RF classifier
Metrics RF SMOTE ADASYN BLS ROS RUS ENN NC S_Tomek CS-RF iCost(Proposed)
Sensitivity 51.10 64.42 64.53 61.20 56.60 84.38 58.18 58.35 64.52 49.22 50.38
Specificity 98.51 96.38 95.94 96.62 97.69 82.78 96.12 96.26 96.35 98.63 98.66
Precision 66.99 64.07 62.41 65.35 67.47 41.70 63.76 64.43 64.08 66.27 67.35
G-mean 63.56 72.93 72.82 69.95 65.36 82.73 64.64 65.01 72.99 61.63 62.60
MCC 54.99 60.28 59.26 59.18 58.04 49.46 56.62 57.10 60.32 53.60 54.76
ROC-AUC 74.81 80.40 80.24 78.91 77.15 83.58 77.15 77.31 80.43 73.92 74.52
F1-Score 55.61 62.80 61.91 61.48 59.43 50.63 58.42 58.82 62.84 54.08 55.18
5.4 Performance in multiclass datasets
The proposed algorithm was also tested in multiclass imbalanced scenarios. A total of 10 datasets with different numbers of classes and different categories (multi-minority, multi-majority, or equal ratio) were utilized in the experiment. We compared the performance of the iCost framework with the traditional CS approach under the OvR setting. Detailed performance measures on individual datasets are provided in the supplementary files. The average performance measures are reported in Table 9. The changes in the MCC score from the proposed algorithm compared to standard CS techniques are illustrated in Fig. 12, Fig. 13, Fig. 14, and Fig. 15 for the LR, SVM, DT, and RF classifiers, respectively.
Multiclass classification is naturally more complicated compared to binary classifications [57]. There are many classes involved and correctly identifying the instances in the minority classes becomes more difficult. Prior to discussing the results, one thing to consider in multiclass scenarios is that even a small change in the MCC or G-mean score is hard to achieve given that there are many classes and the MCC score takes into account predictions in all categories. Achieving higher scores indicates improving correct predictions in minority classes without causing misclassifications of other class instances.
As can be observed from Table 9, the proposed algorithm provides quite a good improvement in overall performance over the standard CS approach. Both the sensitivity and specificity scores improved, indicating the effectiveness of the iCost framework. The precision score along with the F1 and G-mean scores also improved. For the LR classifier, there was almost a 4% improvement in the sensitivity, MCC, and F1 score. Similar improvement was also observed for the SVM classifier. For the DT classifier, the average improvement was very small. However, for none of these cases, there was any decline in the overall scores. Only for the RF classifier, there is a marginal decline in performance. But it is less than 0.05%, making it inconsequential. The drop in performance was mostly from only one dataset, while in other cases, the performance remained almost unchanged.
Table 9: Average performance measures on 10 multiclass datasets
Metrics Accuracy Sensitivity Specificity Precision G-mean MCC F1-Score
LR 82.51 57.86 85.39 69.74 69.28 56.21 59.69
CS-LR 83.22 66.75 88.16 75.20 76.33 61.95 67.88
iCost-LR 84.61 70.27 88.93 77.92 78.56 64.95 71.07
SVM 86.24 68.58 89.44 75.89 77.78 67.31 70.00
CS-SVM 85.70 76.22 92.45 78.28 83.57 70.68 75.63
iCost-SVM 87.39 77.91 93.49 79.87 85.00 73.46 77.76
DT 82.87 75.31 93.44 75.46 83.15 71.25 74.40
CS-DT 82.82 75.37 93.50 75.32 83.24 71.15 74.23
iCost-DT 83.07 75.72 93.65 75.64 83.54 71.73 74.58
RF 87.86 79.58 94.48 81.59 86.26 78.45 80.06
CS-RF 87.98 79.74 94.56 81.48 86.39 78.66 80.03
iCost-RF 87.95 79.67 94.55 81.43 86.35 78.62 79.96
Figure 12: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the LR classifier on 10 multiclass datasets.
Figure 13: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the SVM classifier on 10 multiclass datasets.
Figure 14: Change in MCC score from the iCost algorithm as compared to traditional CS approach for the DT classifier on 10 multiclass datasets.
Figure 15: Change in MCC score from the iCost algorithm as compared to traditional CS approach for the RF classifier on 10 multiclass datasets.
To observe the performance on individual datasets, we can look at Fig. 12, Fig. 13, Fig. 14, and Fig. 15. For the LR classifier, there is a good improvement in performance in most datasets. In the new thyroid dataset, the MCC score improved by around 12%. For the SVM classifier, performance improved significantly in 3 datasets. For others, the improvement was around 1%. For the DT classifier, the performance dropped slightly (0.01%) in 2 datasets. In other cases, the performance improved. For the RF classifier, in the ’balance’ dataset, the MCC score dropped by 1.5%. In most other cases, the changes in the MCC score were very limited. Given the nature of the ensemble algorithms, they provide well-generalized performance. That is why the changes are limited in such an ensemble algorithm.
Overall, our proposed modification in the CSL framework works equally in multiclass classification scenarios as well. Good improvement in performance can be observed, especially in classifiers such as SVM and LR. Ensemble algorithms (such as RF classifiers), which are more robust, are less susceptible to such modifications.
5.5 Performance from MST categorization criteria
Besides the ’neighborhood’ approach, the ’MST’ categorization criterion was also tested on the 65 imbalanced datasets. The detailed performance measures are provided as supplementary files in the GitHub repository. There is an overall improvement in performance over the traditional CS framework from this categorization as well. The changes in the MCC score from the LR classifier are illustrated in Fig. 16. Similar improvements in performance as the ’neighborhood’ approach are noticeable for other classifiers as well.
It is difficult to say which categorization criteria between the two works better. In some datasets, the neighborhood categorization works better, while in others, MST works better. For the LR classifier, in terms of MCC score, the neighborhood categorization gives better performance than the MST categorization in 29 datasets out of 65. For the SVM classifier, however, in only 13 datasets, neighborhood categorization works better. For the DT and RF classifiers, the numbers are 38 and 50, respectively. The difference in performance in many cases is quite negligible.
The difference in the average MCC scores from these approaches on 65 imbalanced datasets is illustrated in Fig. 17. As can be observed from the figure, there is an overall improvement in performance from the proposed approach over the standard CS implementations of the 4 different classification algorithms. There is a slight difference between the two categorization criteria.
Figure 16: Changes in MCC score from the iCost algorithm as compared to traditional CS approach for the LR classifier on 65 imbalanced datasets for MST categorization.
Figure 17: Difference in the average MCC scores between the traditional CS method and the proposed iCost approaches
6 Conclusion
@@
In this study, a modified cost-sensitive learning framework is proposed where data difficulty factors are taken into consideration while penalizing the instances. In the traditional framework, all the instances are weighted equally. This is impractical and can bias the prediction toward the minority-class, leading to an increased amount of false positives. Uniform instance weighting overfits the data by unusually deforming the decision boundary. This does not fare well during testing, resulting in a higher number of misclassifications. To alleviate the issue, we weigh the minority-class samples in the overlapping region more strongly compared to the ones in the non-overlapping region. Carefully assigning higher weights to the hard-to-learn examples while reducing the weights of the other provides a more plausible weighting mechanism, resulting in fewer misclassifications.
We proposed two different instance categorization criteria in this manuscript. We have tested our algorithm on 65 binary and 10 multiclass imbalanced datasets using 4 different classifiers and observed an increment in performance in most cases. The proposed algorithm is computationally light similar to traditional CS approaches. The modification improves the original CS framework implemented in popular Python libraries such as scikit-learn by introducing some reasonable adjustments to it.
@@
Further improvement is possible in the proposed algorithm and we plan to explore those in future works. We measured the instance complexity based on nearest neighbors or MST. However, there are other data complexity measures such as local sets [17] which can be considered. We plan to evaluate the possibility of incorporating those into the framework. While we worked with four different classifiers in this study, the proposed algorithm can be used with other classification algorithms such as XGBoost. The default values for the proposed approach are empirically set. Further research is required in this regard to understand how different data difficulty factors are related to the cost-factor values.
To conclude, this research introduces a novel approach in the cost-sensitive learning domain for imbalanced classification tasks. Penalizing minority-class instances according to their complexity provides improved prediction performance in both binary and multiclass classification scenarios. The implementation details as well as the code are provided to ensure reproducability. The code is written in Python ensuring full compatability with the scikit-learn library. The work presents a new perspective in the field of imbalanced domains and paves the way for the development of new strategies that are based on data-specific characteristics.
7 Appendix
Supplementary files and codes are available in this repository:
https://github.com/newaz-aa/iCost